OBSERVATION OF NUCLEAR SPIN SELECTION RULES IN SUPERSONICALLY EXPANDING PLASMAS CONTAINING H_{3}^{+}

BRIAN A. TOM, MICHAEL B. WICZER, \underline{ANDREW A. MILLS}, KYLE N. CRABTREE, and BEN-JAMIN J. McCALL, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

The reaction of para-H_{2}^{+} with para-H_{2} can form only the para spin modification of H_{3}^{+}, according to nuclear spin selection rules. A proton hop from para-H_{3}^{+} to para-H_{2} maintains the para-H_{3}^{+}, but a hydrogen exchange reaction can produce ortho-H_{3}^{+}. In the hopes of developing a source of H_{3}^{+} that is almost entirely in a single quantum state (J = K = 1), we have investigated a supersonically expanding plasma using a pulsed discharge nozzle with pure para-H_{2} as a precursor gas. The plasma has been interrogated via continous-wave cavity ringdown spectroscopy using a homemade difference frequency laser that combines a tunable Ti:Sapphire laser and a fixed frequency Nd:YAG laser. With this system, we have measured the relative intensities of the R(1, 0), R(1, 1)^{u}, and R(2, 2)^{l} transitions of the ν_{2} band of H_{3}^{+} near 3.67 μm.

We will discuss the construction of the difference frequency laser, our adaptation of continous-wave cavity ringdown spectroscopy to a pulsed source, and the results of our spectroscopic study. We have found that it is possible to produce highly enriched para-H_{3}^{+}, especially when an inert gas is used as a buffer to reduce the rate of reactions between H_{3}^{+} and H_{2}. We have also found that the ratio between the rate coefficients of the proton hop and hydrogen exchange reactions α = k_{H}/k_{E} ∼ 1 at the low temperatures of our expansion, in contrast to the value of α ∼ 2.4 found by the Oka group^{a} at the higher temperatures of a water-cooled hollow cathode discharge.

aM. Cordonnier, D. Uy, R. M. Dickson, K. E. Kerr, Y. Zhang, and T. Oka J. Chem. Phys. \underline{113}, 3181 2000.